OpenStudy (rebeccapink6398):

PLZ CHECK MY WORK! WILL MEDAL!!! A car dealership has 24 sedans and 18 sports coupes on its lot. What is the ratio of sedans to sports coupes on the lot? A. 2 : 3 B. 3 : 4 C. 3 : 2 D. 4 : 3 I KNOW ITS B OR D... plz help

OpenStudy (rebeccapink6398):

@31356 is it b or d?

OpenStudy (anonymous):

It's D. 24/18=4/3

OpenStudy (rebeccapink6398):

Stephanie made 36 muffins. She made 24 apple muffins and the rest were banana. What is the ratio of apple muffins to banana muffins? A. 12 to 36 B. 2 to 1 C. 2 to 3 D. 24 to 36

OpenStudy (rebeccapink6398):

36/24=1.5

OpenStudy (anonymous):

No, it's B. 2 to 1. The banana muffins are 36-24=12. So ratio is 24/12=2/1.

OpenStudy (rebeccapink6398):

oh!

OpenStudy (31356):

Sorry, I wasn't here @RebeccaPink6398. :(

OpenStudy (31356):

You are correct @emcrazy14! :D